# AI, Climate, and Regulation: From Data Centers to the AI Act

### v1

@@ 6 Policy Proposals
Although the AI Act attempts to address climate concerns through various reporting obligations, these measures largely lack consistency and clarity. We identify twelve policy recommendations that should be integrated into the evaluation report due in August 2028 (Article 111(6)), as well as any interpretive guidelines from the AI Office and other agencies, and in reviews and potential textual revisions prior to that date [1]. These measures can be grouped in four categories: Energy and Environmental Reporting Obligations; Legal and Regulatory Clarifications; Transparency and Accountability Mechanisms; and Future Far-Reaching Measures beyond Transparency.
6.1 Energy and Environmental Reporting Obligations
The current AI Act overlooks key environmental factors related to AI systems [13]. The following proposals address this gap:
  • Inclusion of Energy Consumption From Inferences: We propose an interpretation that would at least bring energy consumption for single inferences back into the scope [1]. Adoption of this interpretation by courts, authorities and industry, however, is uncertain. Therefore, both single and overall inferences should be included as a reporting category in Annex XI and Annex XII (vis-à-vis authorities and downstream providers) through delegated acts from the Commission (Articles 53(5) and (6)) and future recommendations from the AI Office.
  • Indirect Emissions and Water Consumption Reporting: The Act currently omits indirect emissions from AI applications (e.g., those used for oil and gas exploration [7]) and water consumption [10]. Reporting should include:
    • Providers reporting water usage, and
    • Deployers reporting application-related emissions, allowing for estimates where precise measurements are impossible, particularly when dealing with future scenarios.
  • Energy Reporting at the Cumulative Server Level: Energy consumption should be reported at the cumulative server level[1]. In this endeavor, estimations may be used only when direct measurements are unavailable. These principles should guide the development of technical standards under Article 40 of the AI Act, as well as the potential implementation of the Global Digital Compact on a global scale.
@@ 6.1 Energy and Environmental Reporting Obligations
Overall, these proposals aim to ensure comprehensive energy and environmental accountability for AI systems, both for development and inference phases.
6.2 Legal and Regulatory Clarifications on AI Models and Providers
@@
Ambiguities in the AI Act regarding provider responsibilities for AI models need to be clarified to ensure that entities know when and what to comply with. The following proposals address these uncertainties:
@@
  • Clarifying Provider Status for Model Modifications: The current definition of provider status needs to focus on substantial modifications to AI models, specifically those that involve changing the model’s weights. Reporting obligations and the change-of-provider threshold should be tied to computational costs incurred during significant modifications, with a minimum computational cost threshold ensuring that only energy-intensive changes trigger additional reporting [1].
@@
  • Elimination of the Open-Source Exemption: The open-source exemption from reporting obligations should be removed, as making parts of a model public does not justify exclusion from environmental accountability [1]. Open-source models can have significant energy implications and should adhere to the same reporting standards as proprietary models.
@@
Overall, these clarifications seek ensure that companies understand their obligations, that open-source models are on par with closed models, and that the regulatory framework sets effective incentives at the source (data centers) to mitigate the environmental impact of AI model development and deployment.
6.3 Transparency and Accountability Mechanisms
@@
  • Public Access to Climate-Related Disclosures: All climate-related disclosures, including energy consumption, should be made accessible to the general public [1]. If only aggregate data are shared, trade secrets can be protected while allowing for public scrutiny by analysts, academics, and NGOs. Public transparency would drive market pressure, reputational incentives, and public accountability; it would, arguably, encourage companies to reduce their environmental impact.
@@
These transparency mechanisms will help foster public accountability and encourage companies to take proactive steps to reduce their environmental footprint.
6.4 Beyond Transparency
@@
  • Restrictions on AI Training and Inference During Peak Hours: One future measure could involve limiting AI training and certain inference tasks (e.g., those for recreation-related purposes) to non-peak energy demand hours [4]. This would reduce the strain on energy grids during peak times and make AI energy usage more manageable and sustainable.
@@
  • Tradable Energy Budgets for AI: Another possible long-term measure could involve implementing tradable energy budgets for AI, similar to the EU Emissions Trading System (EU ETS) [4]. Companies that consume particularly large amounts of energy would have a capped energy budget for AI operations and could trade these credits. This would create a market-based approach to managing and reducing AI-related energy consumption. It would, arguably, encourage companies to optimize their energy usage and invest in energy efficiency.
These forward-looking measures go beyond transparency to address the larger-scale challenges AI poses to energy infrastructure and aim to create sustainable solutions that balance AI’s growth with the global need for clean energy.
@@
Table 1: Summary of Policy Proposals for AI and Environmental Impact
6.5 Interim Conclusion
By implementing our proposals concerning Energy and Environmental Reporting Obligations, Legal and Regulatory Clarifications, Transparency and Accountability Mechanisms, and Future Far-Reaching Measures, the AI Act and related regulation can be significantly strengthened to address both immediate and long-term environmental impacts of AI systems. These measures would create a comprehensive framework that not only ensures compliance with current reporting obligations but also anticipates future energy challenges posed by the growing role of AI. The inclusion of market-based approaches and restrictions on AI operations during peak energy periods may help mitigate the broader environmental impact of AI technologies and help navigate the dual, urgent tranformations of AI and climate change. Table 1 summarizes our proposals.
7 Conclusion
AI systems, many of them containing energy-intensive generative AI components, are increasingly integrated into economic and societal processes. Importantly, as AI intersects with robotics, writing code and steering devices, it is poised to merge decisively into industrial processes, proliferating its deployment, but also its climate effects. Currently, this intersection between the AI in the green transition sits at a regulatory blind spot, inadequately addressed in current regulation, even in highly regulated jurisdictions, such as the EU.
In this paper, we take closer look at sustainability-related data center regulation in the EU and the US, as well as the sustainability provisions in the AI Act. While they present a good first step, they are too often not ambitious enough or present other shortcomings, like a lack of clarity and consistency, or significant challenges in implementation.
To counter these shortcomings, we make twelve concrete proposals, grouped into four areas: the inclusion of energy consumption during inferences, reporting indirect emissions and water consumption, and energy reporting at the cumulative server level under Energy and Environmental Reporting Obligations; clarification of provider status, elimination of the open-source exemption, and renewable energy targets for data centers under Legal and Regulatory Clarifications; public access to climate disclosures, energy reporting for High-Risk AI, and environmental risks in risk assessments under Transparency and Accountability; and restrictions on AI during peak hours, obligations for renewable energy sources, and tradable energy budgets under Future Far-Reaching Measures beyond Transparency.

### v2

@@ 7. Policy Proposals
Although the AI Act attempts to address climate concerns through various reporting obligations, these measures largely lack consistency and clarity. We identify twelve policy recommendations that should be integrated into the evaluation report due in August 2028 (Article 111(6)), as well as any interpretive guidelines from the AI Office and other agencies, and in reviews and potential textual revisions prior to that date (see also (Anonymous, 2024a)). These measures can be grouped in four categories: Energy and Environmental Reporting Obligations; Legal and Regulatory Clarifications; Transparency and Accountability Mechanisms; and Future Far-Reaching Measures beyond Transparency.
7.1. Energy and Environmental Reporting Obligations
@@
The current AI Act overlooks key environmental factors related to AI systems (Luccioni et al., 2024c). The following proposals aim to ensure comprehensive energy and environmental accountability for AI systems, both for development and inference phases.
  • Inclusion of Energy Consumption From Inferences: We propose an interpretation that would at least bring energy consumption for single inferences back into the scope (see also (Anonymous, 2024a)). Adoption of this interpretation by courts, authorities and industry, however, is uncertain. Therefore, both single and overall inferences should be included as a reporting category in Annex XI and Annex XII (vis-à-vis authorities and downstream providers) through delegated acts from the Commission (Articles 53(5) and (6)) and future recommendations from the AI Office.
  • Indirect Emissions and Water Consumption Reporting: The Act currently omits indirect emissions from AI applications (e.g., those used for oil and gas exploration (Kaack et al., 2022)) and water consumption (Li et al., 2023b). Reporting should include: Providers reporting water usage, and deployers reporting application-related emissions, allowing for estimates where precise measurements are impossible, particularly when dealing with future scenarios.
  • Energy Reporting at the Cumulative Server Level: Energy consumption should be reported at the cumulative server level (see also (Anonymous, 2024a)). In this endeavor, estimations may be used only when direct measurements are unavailable. These principles should guide the development of technical standards under Article 40 of the AI Act, as well as the potential implementation of the Global Digital Compact on a global scale.
7.2. Legal and Regulatory Clarifications on AI Models and Providers
@@
Ambiguities in the AI Act regarding provider responsibilities for AI models need to be clarified to ensure that entities know when and what to comply with. The following proposals seek ensure that companies understand their obligations, that open-source models are on par with closed models, and that the regulatory framework sets effective incentives at the source (data centers) to mitigate the environmental impact of AI model development and deployment.
@@
  • Clarifying Provider Status for Model Modifications: The current definition of provider status needs to focus on substantial modifications to AI models, specifically those that involve changing the model’s weights. Reporting obligations and the change-of-provider threshold should be tied to computational costs incurred during significant modifications, with a minimum computational cost threshold ensuring that only energy-intensive changes trigger additional reporting (Anonymous, 2024a).
@@
  • Elimination of the Open-Source Exemption: The open-source exemption from reporting obligations should be removed, as making parts of a model public does not justify exclusion from environmental accountability (Anonymous, 2024a). Open-source models can have significant energy implications and should adhere to the same reporting standards as proprietary models.
@@
7.3. Transparency and Accountability Mechanisms
@@
  • Public Access to Climate-Related Disclosures: All climate-related disclosures, including energy consumption, should be made accessible to the general public (Anonymous, 2024a). If only aggregate data are shared, trade secrets can be protected while allowing for public scrutiny by analysts, academics, and NGOs. Public transparency would drive market pressure, reputational incentives, and public accountability; it would, arguably, encourage companies to reduce their environmental impact.
@@
7.4. Beyond Transparency
@@
  • Restrictions on AI Training and Inference During Peak Hours: One future measure could involve limiting AI training and certain inference tasks (e.g., those for recreation-related purposes) to non-peak energy demand hours (Hacker, 2024b). This would reduce the strain on energy grids during peak times and make AI energy usage more manageable and sustainable.
@@
  • Tradable Energy Budgets for AI: Another possible long-term measure could involve implementing tradable energy budgets for AI, similar to the EU Emissions Trading System (EU ETS) (Hacker, 2024b). Companies that consume particularly large amounts of energy would have a capped energy budget for AI operations and could trade these credits. This would create a market-based approach to managing and reducing AI-related energy consumption. It would, arguably, encourage companies to optimize their energy usage and invest in energy efficiency.
8. Objections and Solutions
Implementing the proposed policy recommendations presents both practical and political challenges, particularly given the EU’s broader regulatory landscape and the evolving international context. This section outlines four key objections to our proposals and sketches solutions. First, a central priority of the new European Commission is the simplification of the digital acquis (‘fitness check’) (MLex, 2025). There is concern that adding further reporting requirements and regulatory obligations could create administrative burdens that counteract this goal. However, several of our proposals, particularly standardized energy reporting metrics and clearer environmental risk assessments, align with the EU’s push for regulatory clarity and streamlining. Rather than adding complexity, these measures would, arguably, make compliance more predictable and feasible.
Second, the shifting international AI governance landscape creates uncertainty regarding the feasibility of ambitious environmental rules. There is concern that EU regulatory efforts might face resistance from international AI providers or risk competitive disadvantages. While acknowledging a trend toward deregulation in some corners of the world, we argue that this reinforces the need for robust EU policy leadership. The AI Act’s extraterritorial scope already compels global AI providers to comply with EU regulations when operating in the European market. By introducing clear, enforceable environmental obligations, the EU can set a global regulatory precedent, and shape international norms despite external pressures. Moreover, some of our proposals - such as market-based mechanisms for energy efficiency - offer flexible compliance pathways that may be more politically acceptable in a deregulatory global environment.
Third, the AI industry, particularly large-scale AI providers and data center operators, may resist additional regulatory obligations, citing economic and technological constraints. There is also concern that lack of enforcement mechanisms could render some obligations ineffective. Hence, our proposals emphasize practical enforcement strategies, including gradual implementation, industry incentives, and transparency-driven accountability. For example: public access to climate-related disclosures can create reputational incentives for compliance, even in the absence of strict enforcement; energy efficiency targets for data centers can be phased in over time for a realistic transition without abrupt regulatory shocks; clarifications on provider status and reporting obligations offer legal certainty to companies. Fourth, some of our more ambitious proposals - such as tradable energy budgets for AI or restrictions on AI training during peak hours - may be viewed as far-reaching or difficult to implement in the near term. While we recognize that not all proposals will be immediately adopted, it is crucial to articulate ambitious policy options at this juncture. The AI regulatory landscape is evolving, and incremental steps toward long-term measures can be integrated into future evaluation rounds of the AI Act, the EED, and related frameworks. As with the EU Emissions Trading System (ETS), initial skepticism toward market-based environmental regulations can give way to gradual acceptance as regulatory mechanisms mature. Given the current political momentum and upcoming AI Act evaluation rounds, there is a narrow but critical window to influence environmental AI governance. While not all recommendations will be immediately adopted, presenting a structured, well-argued policy package ensures that sustainability concerns remain central to AI regulation in the EU and beyond - rather than being sidelined in the face of deregulation.
9. Conclusion
AI systems, many of them containing energy-intensive generative AI components, are increasingly integrated into economic and societal processes. Importantly, as AI intersects with robotics, writing code and steering devices, it is poised to merge decisively into industrial processes, proliferating its deployment, but also its climate effects. Currently, this intersection between the AI and the green transition sits at a regulatory blind spot, inadequately addressed in current regulation. In this paper, we take a closer look at sustainability-related data center regulation and the sustainability provisions in the AI Act. While they present a good first step, they too often lack ambition, clarity and consistency, or present significant challenges in implementation. To counter these shortcomings, we provide interpretations of the AI Act in line with its purpose of environmental protection, provide guidance on operationalizing the requirements, and make twelve concrete proposals, grouped into four areas: Energy and Environmental Reporting Obligations; Legal and Regulatory Clarifications; Transparency and Accountability; and Future Far-Reaching Measures beyond Transparency.
Appendix A Measuring Points for IT Equipment in Data Centers
Figure 1. Measuring points for IT equipment energy consumption, from Annex II of the Delegated Regulation EU/2024/1364 of March 14, 2024
Appendix B Policy Proposals Overview
Table 1. Summary of Policy Proposals for AI and Environmental Impact